what solution set will satisfy if the system of equation 4x+y=8 and 2x=3y+11 are manipulated using substitution?

Answer by rfadrogane(214)


thuswhat solution set will satisfy if the system of equation 4x+y=8 and 2x=3y+11 are manipulated using substitution?
Sol'n
4x+y=8 ---(1)
2x-3y=11 ---(2)
from (1):
y=8-4x ---(1')
substitute (1')into (2)
2x-3(8-4x)=11
2x-24+12x=11
14x=11+24
14x=35
x=35/14
x=5/2
from (1')
y=8-4(5/2)
y=8-10
y=-2
thus, x=5/2 & y=-2 ----answer
^_^